# Every semi-normalized unconditional Schauder frame in Hilbert spaces contains a frame

### v1

@@ 2. Preliminaries
The absolute constant $C$ in Theorem 2.2 may be chosen so that $\sum_{j=0}^{N-1}(B_{j}-1)\,\leq\,C\sqrt{2^{N}\delta},$ for every $N\in\mathbb{N}$ such that $2^{N}\delta<1$ (See [11, Lemma 5.2] or [40, Lemma 10.20] for a proof).
@@
([46, Theorem 1.2]) Let $\{x_{n}\}_{n\in\mathbb{N}}$ be an unconditional Schauder frame for $H$ with coefficient functionals $\{y_{n}\}_{n\in\mathbb{N}}$. Then there exists some sequence of nonzero scalars such that $\{c_{n}\}_{n\in\mathbb{N}}$ such that both $\{c_{n}x_{n}\}_{n\in\mathbb{N}}$ and $\{\bar{c}_{n}^{-1}y_{n}\}$ are frames for $H.\diamondsuit$
@@ 3. Main Results
Let $(m_{n,j})_{j\in\mathbb{N}}\subseteq\mathbb{N}$ be a sequence of scalars that $\sum_{n\in I}\sum_{j=1}^{L_{n}}(2^{-\ell_{n,j}}+2^{-m_{n,j}})=\lceil\sum_{j=1}^{L_{n}}2^{-\ell_{n,j}}\rceil$ and let $\beta\in\mathbb{N}$ be such that $1<2^{\beta}\frac{\epsilon^{2}}{4C^{2}\delta}\leq 2$, where $\lceil\cdot\rceil$ is the ceiling function and $C>0$ is the absolute constant described in Theorem 2.2. We then define $\eta=\max\limits_{n,j}(\ell_{n,j},m_{n,j},\beta)$ and a family of positive trace-class operators $\{\phi_{n}\}_{n\in\mathbb{N}}$ that satisfies
@@
It follow that $\bigl(\,B_{b}\cap\{(n,i)\in I_{1}\,|\,n=k\,\}\,\bigr)\,\leq\,2^{1+\eta-N}c_{k}=2^{\beta+1}c_{k}$ for all binary selectors of order $N$.
@@
Moreover, since $\sum_{(n,i)\in I_{1}}T_{n,i}+\sum_{(n,i)\in I_{2}}\phi_{n,i}=\sum_{n\in I}\sum_{j=1}^{L_{n}}(2^{-\ell_{n,j}}T_{n}+2^{-m_{n,j}}\phi_{n})$, we can even pick $J_{1},J_{2}$ to be the subsets such that
@@
$\frac{\epsilon}{2}P_{M^{\perp}}\,\leq\,P_{M^{\perp}}\bigl(\,\Phi-\Theta\,\bigr)P_{M^{\perp}}\,\leq\,\frac{\epsilon}{2}P_{M^{\perp}},$
@@
It follows that
@@
$(3+\sqrt{2})\sqrt{\gamma}\textbf{I}-\frac{\epsilon}{2}P_{M^{\perp}}\,\leq\,\Phi-T\,\leq\,(3+\sqrt{2})\sqrt{\gamma}\textbf{I}+\frac{\epsilon}{2}P_{M^{\perp}}.$
By the construction of $\phi_{n}$ and the fact that $T_{n,i}=2^{-\eta}T_{n}$, it is equivalent that there exists some sampling function $\sigma\colon I^{\prime}\rightarrow I$ such that
and $\{n\in I^{\prime}\,|\,\sigma(n)=k\}\leq 2^{\beta+1}c_{k}$ for all $k\in I$.
@@ 4. Applications
    $\diamondsuit$
The modulation space $M^{1}(\mathbb{R}^{d})$ is also called Feichtinger algebra in literature. The discovery of modulation spaces can be traced back to early 1980s by Feichtinger. Through a series of subsequent collaborations between Feichtinger and Gröchenig, these spaces have been recognized as fundamental to the development of time–frequency analysis. Loosely speaking, modulation spaces are a family of function spaces which classify functions based on their joint time and frequency localization (via short-time Fourier transform). We remark that, in light of the main focus of this paper, the definition of modulation spaces that is presented above is merely a specific family of modulation spaces commonly found in the literature. For more details and related applications of modulation spaces, we refer to [24] and [8].
@@ 4.1. Olson-Zalik Conjecture
Regarding subspaces of $L^{2}(\mathbb{R}^{d})$ that do not admit Riesz bases, frames or unconditional Schauder frames that are formed by systems of translates, it was shown by the author in [48] that no modulation space $M^{p}(\mathbb{R})$ with $1<p\leq 2$ admits an unconditional Schauder frame of translates with a single generator. Also, no infinite-dimensional closed subspace $S$ of $L^{2}(\mathbb{R}^{d})$ that contains a nonzero function in $M^{1}(\mathbb{R}^{d})$ and is closed under modulation (i.e., there exists some nonzero $b\in\mathbb{R}^{d}$ such that $M_{b}f\in S$ for some nonzero $b\in\mathbb{R}$ and all $f\in S$) admits a frame of translates generated by finitely many functions ([50]). We apply our main result below to characterize a broader class of subspaces that fail to admit unconditional Schauder frames of translates generated by finitely many functions.
@@
By Feichtinger’s conjecture (now known as Feichtinger’s theorem, for example, see [10, Section 4.2]), ${\mathcal{G}}(\phi,\Lambda)$ can be decomposed to into a finite union of subsequences, each of which forms a Riesz basis for its closed linear span. Since ${\overline{\text{span}}}\bigl\{{\mathcal{G}}(\phi,\Lambda)\bigr\}$ is infinite-dimensional, there exists an infinite subset $\Lambda^{\prime}\subseteq\Lambda$ such that ${\mathcal{G}}(\phi,\Lambda^{\prime})$ is a Riesz basis for its closed span. Now by Theorem 3.3, there exist subsets $\Gamma_{j}^{\prime}\subseteq\Gamma_{j}$ such that ${\mathcal{G}}_{N}(g_{j},\Gamma_{j}^{\prime})$ is a frame for $L^{2}(\mathbb{R}^{d}).$ The statement then follows from [50, Corollary 4.4]. ∎
@@
It is known that ${\mathcal{G}}(\phi,\{0\}\times\Lambda)$ is a Bessel sequence when $\Lambda$ is uniformly discrete (for example, see [23, Lemma 3.3]). The statement then follows Theorem 4.2. ∎
@@ 4.2. Beurling density associated with Gabor systems
Consequently, the last missing piece in this direction is to determine the appropriate Beurling density condition for Gabor systems that form unconditional Schauder frames. Due to the absence of the upper frame inequality constraint, it is not hard to see there exist Gabor systems with infinite Beurling density that form unconditional Schauder frames (and hence Schauder frames) if one allows the associated coefficient functionals to contain infinitely many zero functions. The remaining question would be how small the associated Beurling density can be. We resolve this final aspect in the following theorem. Furthermore, we construct unconditional Schauder frames with infinite Beurling density for which all associated coefficient functionals are nonzero. The construction of such unconditional Schauder frames is inspired by [21, Theorem 3.2].
@@ 4.3. Uncertainty Principle
is finite. It was shown by Gröchenig and Malinnikova in [26] that sequences of the form $\{e^{\pi\bar{\lambda}z}\}_{\lambda\in\Lambda}$ is never a Reisz basis for ${\mathcal{F}}^{2}(\mathbb{C}^{d})$ regardless of the choice of $\Lambda\subseteq\mathbb{C}^{d}$. Ascenci et al. then improved this result by showing that $D^{-}(\Lambda)>1$ if $\{e^{\pi\bar{\lambda}z}\}_{\lambda\in\Lambda}$ (see [5, Remark 1.9]) is a frame for ${\mathcal{F}}^{2}(\mathbb{C}^{d})$. Let $\phi(x)=2^{\frac{d}{4}}e^{-\pi\|\,x\,\|^{2}}$ be the Gaussian function in $d$ variables and $\Lambda\subseteq\mathbb{C}^{d}$ be a countable subset. It is known that frames of the form $\{e^{\pi\bar{\lambda}z}\}_{\lambda\in\Lambda}$ (or equivalently, sampling sequences) for ${\mathcal{F}}^{2}(\mathbb{C}^{d})$ are unitarily equivalent to Gabor frames of the form $\{M_{-b}T_{a}\phi\}_{(a+ib)\in\Lambda}$ for $L^{2}(\mathbb{R}^{d})$ via Bargmann transform (for example, see [24, Section 3.4]). Using the Theorem 3.3, Theorem 4.8 and the fact that $\phi\in M^{1}(\mathbb{R}^{d})$, we obtain the following corollary. Here the Beurling density of a countable subset of $\mathbb{C}^{d}$ is defined by the obvious modification of 1.5.
@@ 4.4. Exponential Systems
Combining Theorem 3.3 with [20, Theorem 1.1(i)], we obtain the following theorem. Recall that we say that a sequence of subsets of $\mathbb{R}^{d}$, $\{\Lambda_{n}\}_{n\in\mathbb{N}}$, converges weakly to $\Lambda\subseteq\mathbb{R}^{d}$ if for every $\epsilon>0$ and every $r>0$ there exists $N\in\mathbb{N}$ such that
@@
Assume that $D^{-}(\Lambda)=|S|$. Then by Theorem 3.3 and Landau’s necessary density condition, there exists some subset $\Lambda^{\prime}\subseteq\Lambda$ with $D^{-}(\Lambda)=|S|$ such that ${\mathcal{E}}(\Lambda^{\prime})$ forms a frame for $L^{2}(S).$ The statement then follows by [20, Theorem 1.1(i)]. ∎
Arguing similarly to the one used in [20, Theorem 1.2(i)], we obtain the following result.

### v2

@@ Abstract.
Key words and phrases:
unconditional Schauder frame, frame, Gabor system, Beurling density
2020 Mathematics Subject Classification:
42C15, 42C40
@@ 2. Preliminaries
The absolute constant $C$ in Theorem 2.2 may be chosen so that $\sum_{j=0}^{N-1}(B_{j}-1)\,\leq\,C\sqrt{2^{N}\delta},$ for every $N\in\mathbb{N}$ such that $2^{N}\delta<1$ (See [11, Lemma 5.2] or [42, Lemma 10.20] for a proof).
@@
([48, Theorem 1.2]) Let $\{x_{n}\}_{n\in\mathbb{N}}$ be an unconditional Schauder frame for $H$ with coefficient functionals $\{y_{n}\}_{n\in\mathbb{N}}$. Then there exists some sequence of nonzero scalars such that $\{c_{n}\}_{n\in\mathbb{N}}$ such that both $\{c_{n}x_{n}\}_{n\in\mathbb{N}}$ and $\{\bar{c}_{n}^{-1}y_{n}\}$ are frames for $H.\diamondsuit$
@@ 3. Main Results
Let $\{m_{n,j}\,|\,n\in I,1\leq j\leq M_{n}\}$ be a set of scalars that
$\sum_{j=1}^{L_{n}}2^{-\ell_{n,j}}+\sum_{j=1}^{M_{n}}2^{-m_{n,j}}=\Big\lceil\sum_{j=1}^{L_{n}}2^{-\ell_{n,j}}\Big\rceil\quad\text{for all }n\in I.$
Let $\beta\in\mathbb{N}$ be such that $1<2^{\beta}\frac{\epsilon^{2}}{4C^{2}\delta}\leq 2$, where $\lceil\cdot\rceil$ is the ceiling function and $C>0$ is the absolute constant described in Theorem 2.2. We then define $\eta=\max\limits_{n,j}(\ell_{n,j},m_{n,j},\beta)$ and a family of positive trace-class operators $\{\phi_{n}\}_{n\in\mathbb{N}}$ that satisfies
@@
It follow that
$\bigl(\,B_{b}\cap\{(n,i)\in I_{1}\,|\,n=k\,\}\,\bigr)\,\leq\,2^{1+\eta-N}c_{k}=2^{\beta+1}c_{k}$ (3.1)
for all binary selectors of order $N$.
@@
Moreover, note that the collection $\{B_{b}\,|\,b\in\{0,1\}^{N}\}$ forms a partition of $J_{1}\cup J_{2}$. Consequently, we have
$\begin{split}\sum_{b\in\{0,1\}^{N}}\sum_{(n,i)\in B_{b}}\bigl(\,T_{n,i}+\phi_{n,i}\,\bigr)&\,=\,2^{-\eta}\sum_{b\in\{0,1\}^{N}}\sum_{(n,i)\in B_{b}}\bigl(\,T_{n}+\phi_{n}\,\bigr)\\ &\,=\,\sum_{n\in I}\Bigl(\,\sum_{j=1}^{L_{n}}2^{-\ell_{n,j}}T_{n}+\sum_{j=1}^{M_{n}}2^{-m_{n,j}}\phi_{n}\,\Bigr)\end{split}$ (3.3)
Using the additivity of trace and $\bigl(\,\{0,1\}^{N}\,\bigr)=2^{N}$, we see that there exists at least one binary selector $B_{b}$ such that
$\text{tr}\Bigl(\,2^{N}\sum_{(n,i)\in B_{b}}\bigl(\,T_{n,i}+\phi_{n,i}\,\bigr)\,\Bigr)\,\leq\,\text{tr}\Bigl(\,\sum_{n\in I}\Bigl(\,\sum_{j=1}^{L_{n}}2^{-\ell_{n,j}}T_{n}+\sum_{j=1}^{M_{n}}2^{-m_{n,j}}\phi_{n}\,\Bigr)\,\Bigr)=2\gamma.$
Thus, we can even pick $J_{1},J_{2}$ in Equation (3.2) to be the subsets such that
@@
$\frac{\epsilon}{2}P_{M^{\perp}}\,\leq\,P_{M^{\perp}}\bigl(\,\Phi-\Theta\,\bigr)P_{M^{\perp}}\,\leq\,\frac{\epsilon}{2}P_{M^{\perp}},$ (3.8)
@@
Combining Equation (3.6)–(3.9), it follows that
@@
$(3+\sqrt{2})\sqrt{\gamma}\textbf{I}-\frac{\epsilon}{2}P_{M^{\perp}}\,\leq\,\Phi-\Psi\,\leq\,(3+\sqrt{2})\sqrt{\gamma}\textbf{I}+\frac{\epsilon}{2}P_{M^{\perp}}.$
By the construction of $\phi_{n}$ and the fact that $T_{n,i}=2^{-\eta}T_{n}$, it follows that
$\frac{\epsilon}{2}P_{M^{\perp}}-6\sqrt{\gamma}\textbf{I}\leq\frac{1}{2^{\beta}}\sum_{(n,i)\in J_{1}}T_{n}-T\leq\frac{\epsilon}{2}P_{M^{\perp}}+6\sqrt{\gamma}\textbf{I}.$
Let $I^{\prime}=\{(n,i)\,|\,(n,i)\in J_{1}\}$ and define $\sigma\colon I^{\prime}\rightarrow I$ by $\sigma(n,i)=n.$ It is then equivalent that
$\frac{\epsilon}{2}P_{M^{\perp}}-6\sqrt{\gamma}\textbf{I}\leq\frac{1}{2^{\beta}}\sum_{n\in I^{\prime}}T_{\sigma(n)}-T\leq\frac{\epsilon}{2}P_{M^{\perp}}+6\sqrt{\gamma}\textbf{I}.$
In particular, we have $\{n\in I^{\prime}\,|\,\sigma(n)=k\}\leq 2^{\beta+1}c_{k}$ for all $k\in I$ by Equation (3.1).
We remark the orthogonal decomposition method utilized in the middle of following proof has been used in [24],[11] and [13].
@@ 4. Applications
In particular, the modulation space $M^{1}(\mathbb{R}^{d})$ is usually called Feichtinger algebra in literature.  $\diamondsuit$
The discovery of modulation spaces can be traced back to early 1980s by Feichtinger. Through a series of subsequent collaborations between Feichtinger and Gröchenig, these spaces have been recognized as fundamental to the development of time–frequency analysis. Loosely speaking, modulation spaces are a family of function spaces which classify functions based on their joint time and frequency localization (via short-time Fourier transform). We remark that, in light of the main focus of this paper, the definition of modulation spaces that is presented above is merely a specific family of modulation spaces commonly found in the literature. For more details and related applications of modulation spaces, we refer to [26] and [8].
@@ 4.1. Olson-Zalik Conjecture
Regarding subspaces of $L^{2}(\mathbb{R}^{d})$ that do not admit Riesz bases, frames or unconditional Schauder frames that are formed by systems of translates, it was shown by the author in [50] that no modulation space $M^{p}(\mathbb{R})$ with $1<p\leq 2$ admits an unconditional Schauder frame of translates with a single generator. Assuming that $S$ is an infinite-dimensional closed subspace of $L^{2}(\mathbb{R}^{d})$ that contains a nonzero function in $M^{1}(\mathbb{R}^{d})$ and is closed under modulation (i.e., there exists some nonzero $b\in\mathbb{R}^{d}$ such that $M_{b}f\in S$ for some nonzero $b\in\mathbb{R}$ and all $f\in S$), the author proved in [52] that no such closed subspace admits a frame of translates generated by finitely many functions. We apply our main result below to characterize a broader class of subspaces that fail to admit unconditional Schauder frames of translates generated by finitely many functions.
@@
By Feichtinger’s conjecture (now known as Feichtinger’s theorem, for example, see [10, Section 4.2]), ${\mathcal{G}}(\phi,\Lambda)$ can be decomposed to into a finite union of subsequences, each of which forms a Riesz basis for its closed linear span. Since ${\overline{\text{span}}}\bigl\{{\mathcal{G}}(\phi,\Lambda)\bigr\}$ is infinite-dimensional, there exists an infinite subset $\Lambda^{\prime}\subseteq\Lambda$ such that ${\mathcal{G}}(\phi,\Lambda^{\prime})$ is a Riesz basis for its closed span. Now by Theorem 3.3, there exist subsets $\Gamma_{j}^{\prime}\subseteq\Gamma_{j}$ such that ${\mathcal{G}}_{N}(g_{j},\Gamma_{j}^{\prime})$ is a frame for $L^{2}(\mathbb{R}^{d}).$ The statement then follows from [52, Corollary 4.4]. ∎
@@
It is known that ${\mathcal{G}}(\phi,\{0\}\times\Lambda)$ is a Bessel sequence when $\Lambda$ is uniformly discrete (for example, see [25, Lemma 3.3]). The statement then follows Theorem 4.2. ∎
@@ 4.2. Beurling density associated with Gabor systems
Consequently, the last missing piece in this direction is to determine the appropriate Beurling density condition for Gabor systems that form unconditional Schauder frames. Due to the absence of the upper frame inequality constraint, it is not hard to see there exist Gabor systems with infinite Beurling density that form unconditional Schauder frames (and hence Schauder frames) if one allows the associated coefficient functionals to contain infinitely many zero functions. The remaining question would be how small the associated Beurling density can be. We resolve this final aspect in the following theorem. Furthermore, we construct unconditional Schauder frames with infinite Beurling density for which all associated coefficient functionals are nonzero. The construction of such unconditional Schauder frames is inspired by [22, Theorem 3.2].
@@ 4.3. Uncertainty Principle
is finite. It was shown by Gröchenig and Malinnikova in [28] that sequences of the form $\{e^{\pi\bar{\lambda}z}\}_{\lambda\in\Lambda}$ is never a Reisz basis for ${\mathcal{F}}^{2}(\mathbb{C}^{d})$ regardless of the choice of $\Lambda\subseteq\mathbb{C}^{d}$. Ascenci et al. then improved this result by showing that $D^{-}(\Lambda)>1$ if $\{e^{\pi\bar{\lambda}z}\}_{\lambda\in\Lambda}$ (see [5, Remark 1.9]) is a frame for ${\mathcal{F}}^{2}(\mathbb{C}^{d})$. Let $\phi(x)=2^{\frac{d}{4}}e^{-\pi\|\,x\,\|^{2}}$ be the Gaussian function in $d$ variables and $\Lambda\subseteq\mathbb{C}^{d}$ be a countable subset. It is known that frames of the form $\{e^{\pi\bar{\lambda}z}\}_{\lambda\in\Lambda}$ (or equivalently, sampling sequences) for ${\mathcal{F}}^{2}(\mathbb{C}^{d})$ are unitarily equivalent to Gabor frames of the form $\{M_{-b}T_{a}\phi\}_{(a+ib)\in\Lambda}$ for $L^{2}(\mathbb{R}^{d})$ via Bargmann transform (for example, see [26, Section 3.4]). Using the Theorem 3.3, Theorem 4.8 and the fact that $\phi\in M^{1}(\mathbb{R}^{d})$, we obtain the following corollary. Here the Beurling density of a countable subset of $\mathbb{C}^{d}$ is defined by the obvious modification of 1.5.
@@ 4.4. Exponential Systems
Combining Theorem 3.3 with [21, Theorem 1.1(i)], we obtain the following theorem. Recall that we say that a sequence of subsets of $\mathbb{R}^{d}$, $\{\Lambda_{n}\}_{n\in\mathbb{N}}$, converges weakly to $\Lambda\subseteq\mathbb{R}^{d}$ if for every $\epsilon>0$ and every $r>0$ there exists $N\in\mathbb{N}$ such that
@@
Assume that $D^{-}(\Lambda)=|S|$. Then by Theorem 3.3 and Landau’s necessary density condition, there exists some subset $\Lambda^{\prime}\subseteq\Lambda$ with $D^{-}(\Lambda)=|S|$ such that ${\mathcal{E}}(\Lambda^{\prime})$ forms a frame for $L^{2}(S).$ The statement then follows by [21, Theorem 1.1(i)]. ∎
Arguing similarly to the one used in [21, Theorem 1.2(i)], we obtain the following result.